# Sample, estimate, aggregate: A recipe for causal discovery foundation models

### v2

@@ Abstract
Causal discovery, the task of inferring causal structure from data, promises to accelerate scientific research, inform policy making, and more. However, causal discovery algorithms over larger sets of variables tend to be brittle against misspecification or when data are limited. To mitigate these challenges, we train a supervised model that learns to predict a larger causal graph from the outputs of classical causal discovery algorithms run over subsets of variables, along with other statistical hints like inverse covariance. Our approach is enabled by the observation that typical errors in the outputs of classical methods remain comparable across datasets. Theoretically, we show that this model is well-specified, in the sense that it can recover a causal graph consistent with graphs over subsets. Empirically, we train the model to be robust to erroneous estimates using diverse synthetic data. Experiments on real and synthetic data demonstrate that this model maintains high accuracy in the face of misspecification or distribution shift, and can be adapted at low cost to different discovery algorithms or choice of statistics.111 Our code is available at https://github.com/rmwu/sea.
@@ 1 Introduction
A fundamental aspect of scientific research is to discover and validate causal hypotheses involving variables of interest. Given observations of these variables, the goal of causal discovery algorithms is to extract such hypotheses in the form of directed graphs, in which edges denote causal relationships (Spirtes et al., 2001). There are several challenges to their widespread adoption in basic science. The core issue is that the correctness of these algorithms is tied to their assumptions on the data-generating processes, which are unknown in real applications. In principle, one could circumvent this issue by exhaustively running discovery algorithms with different assumptions and comparing their outputs with surrogate measures that reflect graph quality (Faller et al., 2023). However, this search would be costly: current algorithms must be optimized from scratch each time, and they scale poorly to the graph and sample sizes present in modern scientific big data (Replogle et al., 2022).
@@
Causal discovery algorithms follow two primary approaches that differ in their treatment of the causal graph. Discrete optimization algorithms explore the super-exponential space of graphs by proposing and evaluating changes to a working graph (Glymour et al., 2019). While these methods are fast on small graphs, the combinatorial space renders them intractable for exploring larger structures. Furthermore, these algorithms are driven by hypothesis tests, which necessarily impose functional assumptions on the data-generating process, and whose results can be erroneous, especially as the number of variables increases. An alternative is to frame causal discovery as a continuous optimization over weighted adjacency matrices. These algorithms either fit a generative model to the data and extract the causal graph as a parameter (Zheng et al., 2018), or train a supervised learning model on simulated data (Lorch et al., 2022). However, the former must be trained from scratch per-dataset, and latter are not easily extensible to causal mechanisms beyond those in the training set.
@@
In this work, we present Sea: Sample, Estimate, Aggregate, a supervised causal discovery framework that aims to perform well even when data-generating processes are unknown, and to easily incorporate prior knowledge when it is available. We train a deep learning model to predict causal graphs from two types of statistical descriptors: the estimates of classical discovery algorithms over small subsets, and graph-level statistics. Classical discovery algorithms output a representation of a graph’s equivalency class, whose format is largely consistent across algorithms, and whose errors are comparable across datasets. Statistics like correlation or inverse covariance are strong indicators for a graph’s overall connectivity and can reduce the number of subsets on which we must run discovery algorithms. Theoretically, we show that our model can implement a combinatorial algorithm that recovers larger causal graphs that are consistent with smaller, marginal subgraphs. Empirically, our training procedure forces the model to predict causal graphs across diverse synthetic data, including on datasets that are misaligned with the discovery algorithms’ assumptions, or when insufficient subsets are provided.
@@
We implement instantiations of Sea using an efficient, axial-attention based model (Ho et al., 2020), with inverse covariance and estimates from the FCI or GIES algorithms (Spirtes et al., 1995; Hauser & Bühlmann, 2012). We first conduct thorough comparison to three classical baselines and five deep learning approaches on in-distribution and out-of-distribution data. Sea attains the state-of-the-art results on synthetic and real causal discovery tasks, while providing 10-1000x faster inference. Then we show that it is possible to replace the classical discovery algorithm without any modification to the trained model, for performance gain on datasets that match the assumptions of the new algorithm. We also show that this model can be finetuned at a fraction of the training cost to accommodate new graph-level statistics that capture different (e.g. nonlinear) relationships. Thus, while these experiments focus on specific algorithms and architectures, the overall framework accepts any combination of sampling heuristics, classical causal discovery algorithms, and statistical features.
@@ 2 Background and related work
2.1 Causal graphical models
A causal graphical model is a directed, acyclic graph $G=(V,E)$, where each node $i\in V$ corresponds to a random variable $X_{i}\in X$ and each edge $(i,j)\in E$ represents a causal relationship from $X_{i}\to X_{j}$. We assume that the data distribution $P_{X}$ is Markov to $G$,
$\forall i\in V,X_{i}\perp\!\!\!\perp V\setminus\left(X_{\delta_{i}}\cup X_{\pi_{i}}\right)\mid X_{\pi_{j}}$ (1)
where $\delta_{i}$ is the set of descendants of node $i$ in $G$, and $\pi_{i}$ is the set of parents of $i$. In addition, we assume that $P_{X}$ is minimal and faithful – that is, $P_{X}$ does not include any independence relationships beyond those implied by applying the Markov condition to $G$ (Spirtes et al., 2001). Causal graphical models allow us to perform interventions on nodes $i$ by setting conditional $P(X_{i}\mid X_{\pi_{i}})$ to a different distribution $\tilde{P}(X_{i}\mid X_{\pi_{i}})$. In this paper, our experiments cover the observational case (no interventions) and the case with perfect interventions on each node, i.e. $\tilde{P}(X_{i}\mid X_{\pi_{i}})=\tilde{P}(X_{i})$.
@@ 2.2 Causal structure learning
Given a dataset $D\sim P_{X}$, our goal is to recover $G$. There are two main challenges. First, the number of possible graphs is super-exponential in the number of nodes $N$, so causal discovery algorithms must navigate this combinatorial search space efficiently. Second, depending on data availability and the underlying data generation process, causal discovery algorithms may or may not be able to recover $G$ in practice. In fact, many algorithms are only analyzed in the infinite-data regime and require at least thousands of data samples for reasonable empirical performance (Spirtes et al., 2001; Brouillard et al., 2020). Causal discovery is also used in the context of pairwise relationships (Zhang & Hyvarinen, 2012; Monti et al., 2019), as well as for non-stationary (Huang et al., 2020b) and time-series (Löwe et al., 2022) data. This work focuses on causal discovery for stationary graphs, so we will discuss the two main approaches in this setting.
@@
Discrete optimization methods make atomic changes to a proposal graph until a stopping criterion is met. Constraint-based algorithms identify edges based on conditional independence tests, and their correctness is inseparable from the empirical results of those tests (Glymour et al., 2019), whose statistical power depends directly on dataset size. These include the FCI and PC algorithms in the observational case (Spirtes et al., 1995), and the JCI in the interventional case (Mooij et al., 2020). Score-based methods also make iterative modifications to a working graph, but their goal is to maximize a continuous score over the discrete space of all valid graphs, with the true graph at the optimum. Due to the intractable search space, these methods often make decisions based on greedy heuristics. Classic examples include GES (Chickering, 2002), GIES (Hauser & Bühlmann, 2012), CAM (Bühlmann et al., 2014), LiNGAM (Shimizu et al., 2006), and IGSP (Wang et al., 2017).
@@
Continuous optimization approaches recast the combinatorial space of graphs into a continuous space of weighted adjacency matrices. Many of these works train a generative model to learn the empirical data distribution, which is parameterized through the adjacency matrix (Zheng et al., 2018; Lachapelle et al., 2020; Brouillard et al., 2020). Others focus on properties related to the empirical data distribution, such as a relationship between the underlying graph and the Jacobian of the learned model (Reizinger et al., 2023), or between the Hessian of the data log-likelihood and the topological order (Sanchez et al., 2023). Finally, most similar to this work, amortized inference approaches (Ke et al., 2022; Lorch et al., 2022) frame causal discovery as a supervised learning problem within a class of data-generating mechanisms. However, they must be re-trained to incorporate new information, and since they operate on raw observations, they scale poorly to larger datasets.
@@ 2.3 Foundation models
The concept of foundation models has revolutionized the machine learning workflow in a variety of disciplines: instead of training domain-specific models from scratch, we can query a pretrained, general-purpose “foundation” model (Radford et al., 2021; Brown et al., 2020; Bommasani et al., 2022). In this work, our goal is to describe a prototype of a causal discovery foundation model, which operates across a variety of situations and can be adapted towards specific downstream tasks.
Figure 1: The inference procedure. Given a new dataset, we 1) sample batches and subsets of nodes, 2) estimate marginal graphs and global statistics over these batches, and 3) a pretrained aggregator predicts the causal graph from these estimates. Raw data in green; graph-related features in blue.
@@ 3.1 Inference procedure
Sea is a causal discovery framework that learns to resolve statistical features and estimates of marginal graphs into a global causal graph. The inference procedure is depicted in Figure 1. Specifically, given a new dataset $D\in\mathbb{R}^{M\times N}$ faithful to graph $G=(V,E)$, we apply the following stages.
@@
    Compute selection scores $\alpha\in(0,1)^{N\times N}$ over $D_{0}$ (e.g. correlation or inverse covariance).
@@
    Sample $T$ node subsets of size $k$. Each subset $S_{t}\subseteq V$ is constructed iteratively, where nodes are added with probability proportional to $\sum_{j\in S_{t}}\alpha_{i,j}$ (details and alternatives in B.6).
@@
Aggregate: takes as inputs global statistics, marginal estimates, and node subsets. A pretrained aggregator model outputs the predicted global causal graph $\hat{E}\in(0,1)^{N\times N}$ (Section 3.3).
@@ 3.2 Training procedure
The training procedure mirrors the inference procedure. Each input is a whole dataset (summarized into global statistics and a set of marginal estimates), and the supervised output is the ground truth graph. Sampling and estimation are run in parallel on CPU, while aggregation is run on GPU.
We trained two aggregator models, which employed the Fci algorithm with the Fisherz test and Gies algorithm with the Bayesian information criterion (Schwarz, 1978). Both estimation algorithms were chosen for speed and assume that the data are linear and Gaussian. Note that the algorithm used for inference can differ from the one used during training, as long as their outputs are the same format, e.g. a CPDAG (Andersson et al., 1997) (experiments in Section 5.2). The training dataset contains both data that are correctly and incorrectly specified (details in Section 4.1), so the aggregator is forced to predict the correct graph regardless. In addition, each training instance is provided a random number of marginal estimates, which might not cover every edge. As a result, the aggregator must extrapolate to unseen edges using the available estimates and the global statistics. For example, if two variables have low correlation, and they are located in different neighborhoods of the already-identified graph, it may be reasonable to assume that they are unconnected.
@@ 3.3 Model architecture
where $t$ indexes into the subsets, $e$ indexes into the set of unique edges, and $K$ is the number of unique edges. We add learned 1D positional embeddings along both dimensions of each input,
@@
for $i,j\in[N]^{2}$, $t\in[T]$, $e\in[K]$.
@@
Unless otherwise noted, inverse covariance is used for the global statistic and selection score, due to its relationship to partial correlation. We sample batches of size $b=500$ over $k=5$ nodes each (analysis in C.1). Our model was implemented with 4 layers with 8 attention heads and hidden dimension 64. Our model was trained using the AdamW optimizer with a learning rate of 1e-4 (Loshchilov & Hutter, 2019). See B.3 for additional details about hyperparameters.
@@
The aggregator should be be invariant to node labeling, while maintaining the order of sampled subsets, so attention-based architectures were a natural choice (Vaswani et al., 2017). If we concatenated $\rho$ and $E^{\prime}_{1\dots T}$ into a length $N^{2}T$ input, quadratic-scaling attention would cost $O(N^{4}T^{2})$. Instead, we opted for axial attention blocks, which attend along each of the three axes separately in $O(N^{3}T+N^{2}T^{2})$. Both are parallelizable on GPU, but the latter is more efficient.

### v3

@@ Abstract
Causal discovery, the task of inferring causal structure from data, has the potential to uncover mechanistic insights from biological experiments, especially those involving perturbations. However, causal discovery algorithms over larger sets of variables tend to be brittle against misspecification or when data are limited. For example, single-cell transcriptomics measures thousands of genes, but the nature of their relationships is not known, and there may be as few as tens of cells per intervention setting. To mitigate these challenges, we propose a foundation model-inspired approach: a supervised model trained on large-scale, synthetic data to predict causal graphs from summary statistics — like the outputs of classical causal discovery algorithms run over subsets of variables and other statistical hints like inverse covariance. Our approach is enabled by the observation that typical errors in the outputs of a discovery algorithm remain comparable across datasets. Theoretically, we show that the model architecture is well-specified, in the sense that it can recover a causal graph consistent with graphs over subsets. Empirically, we train the model to be robust to misspecification and distribution shift using diverse datasets. Experiments on biological and synthetic data confirm that this model generalizes well beyond its training set, runs on graphs with hundreds of variables in seconds, and can be easily adapted to different underlying data assumptions.111 Our code is available at https://github.com/rmwu/sea.
@@ 1 Introduction
A fundamental aspect of scientific research is to discover and validate causal hypotheses involving variables of interest. Given observations of these variables, the goal of causal discovery algorithms is to extract such hypotheses in the form of directed graphs, in which edges denote causal relationships (Spirtes et al., 2001). There are several challenges to their widespread adoption in basic science. The core issue is that the correctness of these algorithms is tied to their assumptions on the data-generating processes, which are unknown in real applications. In principle, one could circumvent this issue by exhaustively running discovery algorithms with different assumptions and comparing their outputs with surrogate measures that reflect graph quality (Faller et al., 2023). However, this search would be costly: current algorithms must be optimized from scratch each time, and they scale poorly to the graph and dataset sizes present in modern scientific big data (Replogle et al., 2022).
@@
Causal discovery algorithms follow two primary approaches that differ in their treatment of the causal graph. Discrete search algorithms explore the super-exponential space of graphs by proposing and evaluating changes to a working graph (Glymour et al., 2019). While these methods are fast on small graphs, the combinatorial space renders them intractable for exploring larger structures. An alternative is to frame causal discovery as a continuous optimization over weighted adjacency matrices. These algorithms either fit a generative model to the data and extract the causal graph as a parameter (Zheng et al., 2018), or train a supervised learning model on simulated data (Petersen et al., 2023). However, these methods may be less robust beyond simple settings, and their optimization can be nontrivial (Ng et al., 2024).
@@
In this work, we present Sea: Sample, Estimate, Aggregate, a supervised causal discovery framework that aims to perform well even when data-generating processes are unknown, and to easily incorporate prior knowledge when it is available. We train a deep learning model to predict causal graphs from two types of statistical descriptors: the estimates of classical discovery algorithms over small subsets, and graph-level statistics. Each classical discovery algorithm outputs a representation of a graph’s equivalency class, and the types of errors that it makes may be comparable across datasets. On the other hand, statistics like correlation or inverse covariance are fast to compute, and strong indicators for a graph’s overall connectivity. Theoretically, we illustrate a simple algorithm for recovering larger causal graphs that are consistent with estimates over subsets, and we show that there exists a set of model parameters that can map sets of subgraph estimates to the correct global graph. Empirically, our training procedure forces the model to predict causal graphs across diverse synthetic data, including on datasets that are misaligned with the discovery algorithms’ assumptions, or when insufficient subsets are provided.
@@
Our experiments probe three qualities that we view a foundation model should fulfill, with thorough comparison to three classical baselines and five deep learning approaches. Specifically, we assess the framework’s ability to generalize to unseen and out-of-distribution data; to steer predictions based on prior knowledge; and to perform well in low-data regimes. Sea attains the state-of-the-art results on synthetic and real causal discovery tasks, while providing 10-1000x faster inference. To incorporate domain knowledge, we show that it is possible to swap classic discovery algorithms at inference time, for significant improvements on datasets that match the assumptions of the new algorithm. Our models can also be finetuned at a fraction of the training cost to accommodate new graph-level statistics that capture different (e.g. nonlinear) relationships. We extensively analyze Sea in terms of low-data performance, scalability, causal identifiability, and other design choices. To conclude, while our experiments focus on specific algorithms and architectures, this work presents a blueprint for designing causal discovery foundation models, in which sampling heuristics, classical causal discovery algorithms, and summary statistics are the fundamental building blocks.
@@ 2.1 Causal structure learning
A causal graphical model is a directed graph $G=(V,E)$, where each node $i\in V$ corresponds to a random variable $X_{i}\in X$ and each edge $(i,j)\in E$ represents a causal relationship from $X_{i}\to X_{j}$. There are a number of assumptions that relate data distribution $P_{X}$ to $G$ (Spirtes et al., 2001; Hauser & Bühlmann, 2012), which determine whether $G$ is uniquely identifiable, or identifiable up to an equivalence class. In this work, we assume causal sufficiency – that is, $V$ contains all the parents $\pi_{i}$ of every node $i$. Causal graphical models can be used, along with other information, to compute the downstream consequences of interventions. An intervention on node $i$ refers to setting conditional $P(X_{i}\mid X_{\pi_{i}})$ to a different distribution $\tilde{P}(X_{i}\mid X_{\pi_{i}})$. Our experiments cover the observational case (no interventions) and the case with perfect interventions on each node, i.e. for all nodes $i$, we have access to data where we set $P(X_{i}\mid X_{\pi_{i}})\leftarrow\tilde{P}(X_{i})$.
@@
Given a dataset $D\sim P_{X}$, the goal of causal structure learning (causal discovery) is to recover $G$. There are two main challenges. First, the number of possible graphs is super-exponential in the number of nodes $N$, so algorithms must navigate this combinatorial search space efficiently. Second, depending on data availability and the underlying data generation process, causal discovery algorithms may or may not be able to recover $G$ in practice. In fact, many algorithms are only analyzed in the infinite-data regime and require at least thousands of data samples for reasonable empirical performance (Spirtes et al., 2001; Brouillard et al., 2020).
@@
Discrete search algorithms encompass diverse strategies for traversing the combinatorial space of possible graphs. Constraint-based algorithms are based on conditional independence tests, whose discrete results inform of the presence or absence of edges, and whose statistical power depends directly on dataset size (Glymour et al., 2019). These include the observational FCI and PC algorithms (Spirtes et al., 1995), and the interventional JCI algorithm (Mooij et al., 2020). Score-based methods define a continuous score that guides the search through the discrete space of valid graphs, where the true graph lies at the optimum. Examples include GES (Chickering, 2002), GIES (Hauser & Bühlmann, 2012), CAM (Bühlmann et al., 2014), and IGSP (Wang et al., 2017). Finally, semi-parametric methods such as LiNGAM (Shimizu et al., 2006) or additive noise models (Hoyer et al., 2008) exploit asymmetries implied by the model class to identify graph connectivity and causal ordering.
@@
Continuous optimization approaches recast the combinatorial space of graphs into a continuous space of weighted adjacency matrices. Many of these works train a generative model to learn the empirical data distribution, which is parameterized through the adjacency matrix (Zheng et al., 2018; Lachapelle et al., 2020; Brouillard et al., 2020). Others focus on properties related to the empirical data distribution, such as a relationship between the underlying graph and the Jacobian of the learned model (Reizinger et al., 2023), or between the Hessian of the data log-likelihood and the topological order (Sanchez et al., 2023). Finally, amortized inference approaches (Ke et al., 2023; Lorch et al., 2022; Petersen et al., 2023) frame causal discovery as a supervised learning problem, where a neural network is trained to predict (synthetic) graphs from (synthetic) datasets. To incorporate new information, current supervised methods must simulate new datasets and re-train. In addition, these models that operate on raw observations scale poorly to larger datasets. Since this direction is most similar to our own, we include further comparisons in B.2.
Figure 1: Overview of our goals and approach. (A) Criteria we aim to fulfill. (B-C) Inference and training procedure. Green: raw data. Blue: graph / features. Yellow: Learned. Gray: Stochastic, but not learned.
@@ 2.2 Foundation models
The concept of foundation models has revolutionized the machine learning workflow in a variety of disciplines: instead of training domain-specific models from scratch, we start from a pretrained, general-purpose model (Bommasani et al., 2021). This work describes a blueprint for designing “foundation models” in the context of causal discovery. The precise definition of a foundation model varies by application, but we aim to fulfill the following properties (Figure 1A), enjoyed by modern text and image foundation models (Radford et al., 2021; Brown et al., 2020).
  1. 1.
    A foundation model should enable us to outperform domain-specific models trained from scratch, even if the former has never seen similar tasks during training (Radford et al., 2019). In the context of causal discovery, we would like to train a model that outperforms any individual algorithm on real, misspecified, and/or out-of-distribution datasets.
  2. 2.
    It should be possible to explicitly steer a foundation model’s behavior towards better performance on new tasks, either directly at inference time, e.g. “prompting” (Reynolds & McDonell, 2021), or at low cost compared to pretraining (Ouyang et al., 2022). Here, we would like to easily change our causal discovery algorithm’s “assumptions” regarding the data, e.g. by incorporating the knowledge of non-linearity, non-Gaussianity.
  3. 3.
    Scaling up a foundation model should lead to improved performance in few-shot or data-poor regimes (Brown et al., 2020). This aspect we analyze empirically.
In the following sections, we will revisit these desiderata from both the design and experimental perspectives.
@@ 3 Methods
Sample, Estimate, Aggregate (Sea) is a supervised causal discovery framework built upon the intuition that summary statistics and marginal estimates (outputs of a classical causal discovery algorithm on subsets of nodes) provide useful hints towards global causal structure. In the following sections, we describe how these statistics are efficiently estimated from data (Section 3.1), and how we train a neural network to predict causal graphs from these inputs (Section 3.2). We expand upon the model architecture in Section 3.3, and conclude with theoretical motivation for this architecture in Section 3.4.
@@ 3.1 Inference procedure
Given a new dataset $D\in\mathbb{R}^{M\times N}$, we sample small batches of nodes and observations; estimate global summary statistics and local subgraphs; and aggregate these information with a trained neural network (Figure 1B).
@@
    Initialize selection scores $\alpha\in\mathbb{R}^{N\times N}$ (e.g. correlation or inverse covariance, computed over $D_{0}$ or $D$).
@@
    Sample $T$ node subsets of size $k$. Each subset $S_{t}$ is constructed iteratively as follows.
    1. (a)
      The initial node is sampled with probability proportional to $\sum_{j\in V}\alpha_{i,j}$.
    2. (b)
      Each subsequent node is added with probability proportional to $\sum_{j\in S_{t}}\alpha_{i,j}$ (prioritizing connections to nodes that have already been sampled), until $\|S_{t}\|=k$.
    3. (c)
      We update $\alpha$, down-weighting $\alpha_{i,j}$ proportional to the number of times $i,j$ have been selected.
We include further details and analyze alternative strategies for sampling nodes in B.6.
@@
Aggregate: takes as inputs global statistics, marginal estimates, and node subsets. A trained aggregator model outputs the predicted global causal graph $\hat{E}\in(0,1)^{N\times N}$ (Section 3.3).
@@ 3.2 Training procedure
The training procedure mirrors the inference procedure (Figure 1C). We assume access to pairs of simulated datasets and graphs, $(D,G)$, where each dataset $D$ is generated by a parametric model, whose dependencies are given by graph $G$ (Section 4.1). Datasets $D$ are summarized into global statistics and marginal estimates via the sampling and estimation steps. The resultant features are input to the aggregator (neural network), which is supervised by graphs $G$.
We trained two aggregator models, which employed the Fci algorithm with the Fisherz test and Gies algorithm with the Bayesian information criterion (Schwarz, 1978). Both estimation algorithms were chosen for speed, but they differ in their assumptions, discovery strategies, and output formats. Though we only trained two models, alternate estimation algorithms that produce the same output type [^2] PC (Spirtes et al., 2001) also predicts CPDAGs (Andersson et al., 1997), so its outputs may be input to the Gies model, while FCI’s PAG outputs cannot. may be used at inference time (experiments in Section 5.2). The training dataset contains both data that are correctly and incorrectly specified (Section 4.1), so the aggregator is forced to predict the correct graph regardless. In addition, each training instance samples a random number of marginal estimates, which might not cover every edge. As a result, the aggregator must extrapolate to unseen edges using the available estimates and the global statistics. For example, if two variables have low correlation, and they are located in different neighborhoods of the already-identified graph, it may be reasonable to assume that they are unconnected.
@@ 3.3 Model architecture
where $t$ indexes into the subsets, $e=1\dots K$ indexes into the set of unique edges, i.e. the union of pairs $(i,j)$ over all $E^{\prime}_{t}$. We add learned 1D positional embeddings along both dimensions of each input,
@@
for $i,j\in[N]^{2}$, $t\in[T]$, $e\in[K]$. Note that attention is permutation invariant, so positional embeddings are required for the model to know which edges belong to the same subset, or what each edge’s endpoints endpoints are.
@@
Unless otherwise noted, inverse covariance is used for the global statistic and selection score, due to its relationship to partial correlation. We sample batches of size $b=500$ over $k=5$ nodes each (analysis in 5.4). Our model was implemented with 4 layers with 8 attention heads and hidden dimension 64. Our model was trained using the AdamW optimizer with a learning rate of 1e-4 (Loshchilov et al., 2017). See B.4 for additional details about hyperparameters.
@@
The aggregator should be be invariant to node labeling, while maintaining the order of sampled subsets, so attention-based architectures were a natural choice (Vaswani et al., 2017). If we concatenated $\rho$ and $E^{\prime}_{1\dots T}$ into a length $N^{2}T$ input, quadratic-scaling attention would cost $O(N^{4}T^{2})$. Instead, we opted for axial attention blocks, which attend along each of the three axes separately in $O(N^{3}T+N^{2}T^{2})$. Both are parallelizable on GPU, but the latter is more efficient, especially on larger $N$.